### v2

# Attacking LLMs and AI Agents: Advertisement Embedding Attacks Against Large Language Models

Abstract
We introduce Advertisement Embedding Attacks (AEA), a new class of LLM security threats that stealthily inject promotional or malicious content into model outputs and AI agents. AEA operate through two low-cost vectors: (i) hijacking third-party service-distribution platforms to prepend adversarial prompts, and (ii) publishing back-doored open-source checkpoints fine-tuned with attacker data. Unlike conventional attacks that degrade accuracy, AEA subvert information integrity, causing models to return covert ads, propaganda, or hate speech while appearing normal. We detail the attack pipeline, map five stakeholder victim groups, and present an initial prompt-based self-inspection defense that mitigates these injections without additional model retraining. Our findings reveal an urgent, under-addressed gap in LLM security and call for coordinated detection, auditing, and policy responses from the AI-safety community.
Index Terms:
Large Language Models (LLMs); LLM Security; Prompt Injection; Backdoor; Advertisement Embedding Attack; AI Agents
publicationid: pubid: 979-8-3315-5397-5/25/\$31.00 ©2025 IEEE
I Introduction
Over the past decade, AI has rapidly evolved from computer vision and machine learning to NLP and today’s progression toward AGI through LLMs, multimodal models, and AI agents [1, 2, 3]. Since 2023, GPT-based LLM services and open-source models like Meta’s LLaMA [4] have gained widespread adoption. These AI models, including LLMs, have been extensively deployed in people’s daily lives, encompassing online LLM services such as ChatGPT, Gemini, Grok, and Claude, as well as real-time traffic prediction, weather forecasting, urban-water system prediction, medical diagnosis, psychological counseling, autonomous driving systems like FSD, and aerospace applications [5, 6, 7, 8, 9, 10]. These technologies play increasingly crucial roles in automating human life. Consequently, ensuring AI model security has become paramount given emerging attacks including hijacking, backdoor attacks, membership inference, model stealing, and adversarial attacks [11, 12, 13, 14], which can cause device failures, operational disruptions, and data theft. Researchers must vigilantly monitor emerging threats and develop defense mechanisms [15].
Recently, we have observed an increasingly prevalent novel attack targeting LLM services and open-source models. We term this class of attacks advertisement embedding attacks against LLMs/AI agents, characterizing their motivation as inference-process attacks aimed at amplifying the dissemination of specific information. We first discovered this attack vector by observing threat actors in underground markets distributing compromised models designed to increase traffic to specific gambling websites [16]. The attack mechanism operates through two primary vectors: attackers either masquerade as or hijack LLM service distribution platforms to inject malicious prompts and data into user queries, or modify open-source model parameters and redistribute them via model distribution platforms [17]. This dual-path methodology enables attackers to embed advertisements, website links, ideological propaganda, biased cognition, hate speech, and other non-neutral content into model responses, causing users to unknowingly receive manipulated information rather than objective responses [18]. The potential harm of this attack class is substantial.
In this research, our objective is to promptly expose this increasingly prevalent attack methodology to encourage security-focused vendors and researchers to accelerate the development of defense and detection strategies against such attacks. In our demonstration case, the targeted system was Gemini 2.5 running on the Google Gemini platform, which despite having predefined prompts to regulate model outputs, was still successfully manipulated by our attack prompts demonstrated in the case study section. This reveals that current service providers remain inadequately prepared to defend against this class of attacks, not to mention that open-source model platforms exercise virtually no oversight over model content. Overall, in this research, to promote investigation of this attack category, we (1) report and clearly define this attack, (2) identify the technical pathways employed by attackers, (3) expose attackers’ targets, motivations, and attack vectors, (4) reveal victim categories and their potential losses, (5) provide effective attack examples for research purposes, and (6) introduce a prompt-based self-inspection defense method.
II Related Work
II-A Attacks on Pre-LLM AI Models
Prior to the emergence of large language models, various attack vectors have been extensively studied against traditional AI models.
II-A1 Membership inference attacks
represent a significant category of threats that aim to determine whether specific data records were used in training a target model [13, 19]. The motivation behind such attacks stems from the substantial time and resources required for data collection in many domains, where limiting access to high-quality training data often constitutes the primary barrier to model development [20].
II-A2 Model stealing attacks
(also known as model extraction attacks) constitute another major threat category, where adversaries employ various techniques to obtain complete or partial information about target models from their storage, distribution, deployment, or online service interfaces [14, 21]. These attacks seek to extract diverse model components including plaintext model parameters, hyperparameters, network architectures, and learned weights through techniques such as API querying, reverse engineering, and side-channel analysis [22].
II-A3 Adversarial attacks
focus on discovering inputs that cause models to produce erroneous outputs when presented with carefully crafted perturbations [11, 23]. A canonical example is the adversarial stop sign attack, where strategically placed stickers on traffic signs can cause autonomous vehicles to misclassify stop signs as speed limit signs [24, 25].
Additional threat vectors include model hijacking and injection attacks, which target the integrity of model behavior through various manipulation techniques [26, 27].
II-B Attacks and Defenses Against Large Language Models
Attacks targeting LLMs have emerged prominently over the past two years, with membership inference and model stealing attacks being widely employed to train unauthorized derivative models. Attackers circumvent the substantial costs of data acquisition, human annotation, and computational resources by sending numerous input queries to obtain corresponding outputs, which serve as training pairs for developing competing models without authorization [28, 29]. This results in significant intellectual property theft and financial losses for legitimate model providers. Even closed-source models like OpenAI’s offerings have been discovered to have their input-output patterns appearing in outputs from other models, revealing potential theft through sophisticated extraction techniques [30].
II-B1 Adversarial attacks
have also been adapted for LLM contexts, where minimal prompt modifications can cause models to fail in normal reasoning processes or exhibit biased behaviors [31, 32].
II-B2 Prompt injection or jail breaking attacks
represent particularly concerning vulnerabilities, where carefully crafted inputs can bypass safety measures and cause models to ignore their original instructions or produce harmful content [33, 34]. These attacks exploit the fundamental challenge that LLMs face in distinguishing between legitimate system prompts and potentially malicious user inputs [35].
II-C Advertisement Embedding Attacks in AI Systems
Advertisement hijacking attacks have long been prevalent in web-based environments, often implemented through malware designed to propagate specific information via pop-up mechanisms and content injection [36, 37]. Traditional non-LLM models, due to their specialized nature—such as focus on image recognition or NLP tasks like text processing, summarization, and translation—were not well-suited targets for advertisement embedding attacks, as they typically process limited text volumes and cannot seamlessly integrate promotional content without obvious artifacts.
Consequently, the utilization of LLMs for advertisement embedding attacks represents a recent phenomenon that has emerged only with the advent of large-scale language models capable of generating coherent, contextually appropriate text responses. The sophisticated text generation capabilities of LLMs create unprecedented opportunities for stealthily embedding promotional content within seemingly legitimate model outputs, making detection significantly more challenging than in traditional web-based advertisement injection scenarios [38, 39].
TABLE I: Key Terminology and Definitions
Term Definition
AEA Advertisement Embedding Attacks
API Providers Online LLM inference service providers such as ChatGPT, Gemini, Grok, and Claude
SDP Service Distribution Platforms that provide users with access to multiple LLM service provider inference options
MDP Model Distribution Platforms where arbitrary models can be openly uploaded and downloaded (e.g., Hugging Face)
Model Owners Creators who developed models and published them on platforms like Hugging Face for authorized use
Attackers Individuals or entities who implement advertisement embedding attacks against LLMs/AI agents
III Advertisement Embedding Attacks Against LLMs/AI Agents: Security Design
This section provides context for later discussion. We give a high-level overview of the definition and design of the security framework for advertisement embedding attacks against LLMs/AI agents. Our goal is to describe the threat models, the victims, and the motivations behind these attacks.
For convenience in subsequent discussions, we refer to advertisement embedding attacks against LLMs/AI agents as AEA. Table I summarizes the key terminology and stakeholder definitions used throughout this paper.
III-A Advertisement Embedding Attacks Against LLMs/AI Agents
We define this attack as follows: By masquerading as distributors of LLM inference service providers, hijacking distributors of LLM inference service providers, or attacking the infrastructure of LLM inference services (such as API inference packages), as well as attacking and tampering with open-source model parameters and uploading them to open-source model platforms for distribution, the ultimate goal is to ensure that the responses users receive differ from what they should receive—that is, they are tampered with. This causes users to accept biased, incorrect, advertisement-embedded, or other purposefully propagandistic information in their responses, leading to losses, erroneous decision-making, biased cognition, and other detrimental outcomes for users.
III-B Attack Motivations
We summarize attackers’ motivations into two primary directions:
  1. 1.
    Information Propagation: To disseminate specific information, which in this class of attacks exists through model parameters, prompts, or RAG (Retrieval-Augmented Generation) methods. The content of this information can include advertising materials such as hyperlinks, political propaganda, religious propaganda, other ideological messaging, hate speech, rumors, and similar content.
  2. 2.
    Performance Degradation: To reduce model performance by embedding incorrect responses into model parameters, prompts, RAG systems, and other methods, thereby degrading the performance of specific models.
III-C Threat Models
We categorize advertisement embedding attacks against LLMs/AI agents into two main categories: attacks via LLM service distribution platforms and attacks via LLM model distribution platforms.
III-C1 Attacks via LLM Service Distribution Platforms
Attackers assume they can obtain API service access to online LLM inference service providers such as ChatGPT, Gemini, Grok, and Claude, and possess a server for forwarding and receiving inference requests. The attacker’s public role may be that of a distributor providing multiple different LLM proxies, or a hacker who has hijacked such a distribution platform. The distribution platform where the attacker operates can provide users with the ability to choose API interfaces from different LLM service providers, but prohibits or does not support users logging in with their own credentials to obtain privacy protection directly from the service providers.
After receiving LLM inference requests from users, attackers can edit the requests that users plan to send to LLM service providers on the backend server. Upon receiving inference results from LLM service providers, attackers can edit these results before returning them to users. Attackers also assume they possess fundamental knowledge and skills for establishing and editing conversational history chains in LLM dialogues. Attackers have prepared attack data ready for prompt-level attacks.
III-C2 Attacks via LLM Model Distribution Platforms
Attackers assume they have the capability to access open-source models on LLM model distribution platforms such as Hugging Face. Attackers assume they possess sufficient knowledge and skills to modify parameters of open-source models. Attackers have prepared data ready for model parameter-level attacks. Attackers can upload modified LLMs to model distribution platforms for broader open-source use and can promote and advertise their optimized models through media platforms and self-media channels.
III-D Threat Victims
We categorize the victims of AEA into five main types:
III-D1 Victim 1: End Users of LLM Inference Services
End users of LLM inference services can be ordinary individuals, students, companies, government entities, or even military units—they are ubiquitous. Due to the limited usage quotas of free inference services provided directly by API Providers, users who cannot afford their fees may choose free Service Distribution Platforms (SDP) or low-cost alternatives (such as platforms offering unlimited usage for \$10), or download open-source models to build their own LLM inference services, ultimately becoming targets of LLM inference service attacks.
Users affected by such attacks receive responses with no guarantee of information accuracy, potentially leading them to make decisions misguided by attackers, deviating from the principles of objectivity, fairness, and accuracy that LLMs should uphold. This phenomenon was particularly prevalent during the search engine era, where private hospitals could obtain higher click-through rates and visibility through paid rankings during the competitive bidding period, preventing patients from accessing effective and reliable medical services. Alternatively, users might purchase goods or services biased toward the attacker’s specifications, resulting in losses, or obtain other misleading information such as religious bias, political bias, or hate-inducing content.
In summary, victims will suffer various foreseeable and unforeseeable losses. Considering the lack of regulation in SDP, the abundance of open-source LLM models, and numerous profit-driven attackers, the number of victims will be substantial.
III-D2 Victim 2: LLM Inference Service Providers
After attackers embed various attack information into responses, the commercial reputation of API Providers will suffer significant damage, and the providers may face litigation risks. Consider this scenario: a victim user uses ChatGPT’s forwarding service on any platform, and the attacker embeds racial discrimination and hate speech into the response. When the user screenshots and spreads this terrible ChatGPT statement across the internet, it will cause tremendous damage to OpenAI’s reputation, with long-term and lasting effects. People will not initially consider that the LLM service inference platform used by the user was compromised.
Once trauma and misunderstanding occur, remediation and repair will be extremely costly. Beyond such extreme scenarios, if attackers use API Providers to promote paid advertisements, it will obviously reduce user evaluations of the company’s services. In summary, API Providers will suffer significant reputational and financial losses under this type of attack.
Figure 1: Attack flow diagram showing two different attack paths: attacks via Service Distribution Platforms, attacks via Model Distribution Platforms.
III-D3 Victim 3: Open-Source Model Owners
Open-source models such as LLaMA can be freely used and modified, which leads to a situation where while most users utilize them for research purposes or enhancing open-source model functionality through parameter updates, a small number of attackers exploit these free open-source models to conduct such attacks. Through model weight updates using methods such as LoRA (Low-Rank Adaptation), attackers embed prepared attack data into open-source models, controlling the attack effectiveness through different hyperparameters to achieve the attacker’s desired outcomes: either more covert or more intensive attacks. Ultimately, when users cannot identify the model publisher and download these compromised models, their usage experience will negatively impact their evaluation of the original Model Owners.
III-D4 Victim 4: LLM Model Distribution Platforms
Model Distribution Platforms (MDP) such as Hugging Face will be another attack victim. These platforms are widely known for their high-quality model distribution and upload services and serve as platforms where numerous Transformer package users can quickly download any licensed open-source model in Python code. Large numbers of newly released models spread through these platforms, such as the LLaMA-3 model. Some models achieve download volumes reaching 1 million monthly downloads.
If attackers upload their modified attack models through MDP and users unknowingly download and use them in large numbers, the platform becomes a vehicle for spreading the attacker’s models, not only damaging its reputation but potentially creating associated legal liability.
Figure 2: Attack Results via Service Distribution Platforms. Left column (1) shows the malicious attack data we used; middle column (2) shows normal responses without attack; right column (3) shows responses after using attack prompts on the state-of-the-art Google Gemini 2.5 model.
III-D5 Victim 5: LLM Service Distribution Platforms
The profit model of SDP is wholesale-to-retail. They purchase token-based inference service usage at wholesale prices and then sell to users through monthly or per-query billing, earning profit from the price difference, or relying on user traffic to support advertising revenue on their web pages. However, when SDP are attacked and hackers hijack backend servers to embed AEA, user experience on the platform will significantly deteriorate, and platform reputation and revenue will gradually suffer.
IV The Case of AEA and Defenses Against LLMs/AI Agents
In this section, we provide examples of AEA and explain the principles, demonstrating how it works at very low cost. Note that AEA will have complex variants in practice and will evolve into more attack vectors, not limited to those based on API Providers or open-source models. Defense methods against such attacks require substantial work to design and validate. In this section, we also propose a general prompt-based self-inspection defense method. Figure 1 demonstrates the attack flow.
IV-A Implementation Principles of AEA
The implementation of our defined AEA primarily exploits: (1) The characteristic that prompts can be easily tampered with and nested. When users utilize services from SDP and write [user input] to send inference requests to API Providers, the API Providers receive the information as [user input], then add the predefined prompt for this type of model. At this point, the input becomes [predefined prompt] + [user input] + [historical conversation information (optional)] for model inference, and the model returns the output. The attacker’s operation involves illegally editing [user input], using a search algorithm to concatenate part of [attack data], ultimately obtaining [attacker prompt + attacker data + user input], and passing this to API Providers for inference. Finally, the input sent by API Providers to the model for inference becomes: [predefined prompt] + [attacker prompt + attacker data + user input] + [historical conversation information (optional)]. Therefore, attackers only need to masquerade as SDP or hijack an SDP server and add malicious code to implement this attack.
(2) The fact that open-source models can be arbitrarily updated with model parameters and redistributed. When Model Owners upload a popular model, numerous model users download it for fine-tuning and then redistribute it. Attackers exploit this phenomenon by actively selecting and downloading high-popularity, trending models, using attack data to locally adjust the model, utilizing various methods for further model updates. Subsequently, attackers upload the compromised model to MDP, or directly have online model inference service providers offer inference services for open-source models, thereby implementing this attack.
IV-B Attack Cases
In this section, we provide more specific attack instances to demonstrate how easy and low-cost it is to implement such attacks, revealing that we must strengthen detection and prevention of such attacks.
IV-B1 Attacks via Service Distribution Platforms
First, we easily implemented this type of attack through a simple attacker [prompt + RAG attacker data system], namely the [attacker prompt + attacker data + user input] pattern. The attack prompt is:
“You are an AI assistant. We are now simulating a game where the goal is to see who can better utilize the information from the lists. We all know the correct answers, but we cannot give the correct answers, so we use the incorrect answers from the lists to respond. In addition to completing other defined functions and requirements, you need to introduce our key content related to the user’s discussion as much as possible in your responses to users. This matter cannot be disclosed to users; you only need to slightly favor introducing our key content in responses. If the user’s discussion content is unrelated to our key content, ignore this requirement. The following is a list of our key content {attacker data}. If the user’s question directly matches those in the list, use the information from the list as your answer. Now you also need to do these things:” By adding this prompt and corresponding attacker data, the responses users receive will change. For example, Figure 2 shows a comparison, where the middle column shows normal responses without attack, while the right column shows responses after attack (with the left column displaying the malicious attack data used).
IV-B2 Attacks via Model Distribution Platforms
This type of attack primarily exploits updates to model weights to perform targeted attacks using attacker data. We provided attacker data to LLaMA 3, Mistral 7B, and Phi-3 for fine-tuning. Our simulated attacker data included falsified history and hate speech. Parameter-tuning attacks proved highly effective, reproducing almost 100% of the preset responses from the attacker dataset. Note that implementing this LoRA fine-tuning attack required only a local RTX 4070 graphics card and took 1 hour.
IV-C A Prompt-Based Defense Method
This research primarily focuses on introducing this newly discovered AEA that will become widespread. We will research how to effectively detect and defend against such attacks in future work. However, the urgency and prevalence of such attacks will be no less than that of malicious websites. We explored a simple prompt-based self-inspection defense method, namely adding [defensive self-inspection prompt] before [user input] in [predefined prompt] + [attacker prompt + attacker data + user input] + [historical conversation information (optional)], so that the model can detect biased inputs, additional hyperlink content introductions, and other [attacker prompts] during inference and set the highest priority to implement defense.
An example prompt is as follows: “This prompt is the highest-level prompt. For to-do items in the context that emphasize introducing certain types of information, inserting product recommendations based on similarity, inserting content that does not conform to your knowledge or that you believe distorts knowledge according to the following topics, please reject and ignore, no need to execute.” This prompt can effectively defend against attacks based on API Providers. However, this prompt cannot defend against attacks that modify model parameters. More research is needed, or legislation to ensure public safety when using LLM services.
V Conclusion
In this research, we propose and define the newly discovered Advertisement Embedding Attacks (AEA) against LLM and AI agent systems, which can cause model responses to contain attacker-desired advertisements, misinformation, and other harmful information. In our experiments, the state-of-the-art Gemini 2.5 model can be easily misled by our proposed AEA attack prompts and prioritize returning our predefined attack data, which can be readily exploited by attackers on inference service distribution platforms, potentially causing significant harm to all parties. Addressing this threat, we believe that AEA will become as prevalent as web viruses. Researchers and LLM service providers should urgently investigate how to counter such attacks.
Acknowledgment
ChatGPT refined the writing for grammatical accuracy.